The boiling point of amphetamine, C_9H_13N, is 195
Answer #1

ln(P2/p1) = \DeltaH/R(1/T1-1/T2)

ln(1.733/101.3) = \DeltaH/R(1/468-1/368)

-4.068             = \DeltaH/R(0.00213-0.00271)

-4.068             = \DeltaH/R(-0.00058)

\DeltaH/R           = 7013.79

ln(P/101.3)     = 7013.79(1/468 - 1/293)

                   = 7013.79(0.00213 -0.00341)

                  = 7013.79*-0.00128

                  = -8.977

p/101.3      = 10-3.89

    P           = 1.29*10-4 *101.3

                 = 0.013 torr

PV = nRT

n/V = P/RT

   P = 0.013/760 = 1.71*10-5 atm

n/V   = 1.71*10-5 /0.0821*293

      = 7.1*10-7 mole/L

    = 7.1*10-7 *1000L/m3 *135 = 0.0958 g/m3